OpenStudy (anonymous):

An automobile travels on a straight road for 34 km at 42 km/h. It then continues in the same direction for another 34 km at 84 km/h. (a) What is the average velocity (in km/h) of the car during this 68 km trip? (Assume that it moves in the positive x direction.) (b) What is the average speed (in km/h)?

OpenStudy (anonymous):

a) time for 1st stretch, t1=34/42 hrs time for 2nd stretch, t2=34/84 hrs total time,t=t1+t2 avg. vel.=total displacement/total time

OpenStudy (anonymous):

b) Here avg.speed=avg. vel because displacement and distance traveled are the same